OpenStudy (anonymous):

simplify the number using the imaginary unit i square root - 28

OpenStudy (anonymous):

\[\sqrt{-28}=\sqrt{28}.\sqrt{-1}=\sqrt{4}\sqrt{7} (i)=2\sqrt{7}i\]

OpenStudy (anonymous):

@MtMtx thanks!! but thats not a choice

OpenStudy (anonymous):

it is. Its the last one

OpenStudy (anonymous):

my answer is the same as the last one on your list\[2i \sqrt{7}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

hey @myko <3 ;) hehe

OpenStudy (anonymous):

:)